OpenStudy (anonymous):

Solve 3y+4y< or equal to -35

OpenStudy (anonymous):

start with \[7y\leq -35\] then finish by dividing by 7

OpenStudy (anonymous):

-5?

OpenStudy (anonymous):

yes \[y\leq -5\]

OpenStudy (anonymous):

so it would be {y|y> or equal to -5?

OpenStudy (anonymous):

do we switch the sign when it is negative?

OpenStudy (anonymous):

it is just as i wrote, \[\{y|y\leq -5\}\] you only change the inequality when multiplying or dividing by a negative number, and 7 is positive, so leave it as it started